Calculate the pH of the following solutions.

Solution prepared by diluting 44.0 mL of 6.5×10−2 M Ba(OH)2 to a volume of 323.5 mL .
Answer #1

M1V1 = M2V2

M1 = 6.5*10^-2 M

V1 = 44.0 ml

M2 = ?

V2 = 323.5 ml

6.5*10^-2 *44 = M2*323.5

M2 = 6.5*10^-2 *44/323.5

M2 = 8.84*10^-3 M

[OH-] = M2 = 8.84*10^-3

Ph = 14 - Poh

Poh = -log[OH-]

Poh = -log[8.84*10^-3] = 2.05

Ph = 14 -2.05 = 11.95

or

Ph of Ba(OH)2 is 12